1 Selected Publications (plus my Ph.D. thesis)

[1] D. Avis, D. Bremner, H.xR. Tiwary, and O. Watanabe. Polynomial size linear programs for problems in P. Discrete Applied Mathematics, March 2018. Accepted, with minor revisions.

https://arxiv.org/abs/1408.0807  .

[2]  ,  O.  Devillers,

M. Glisse, S. Lazard, G. Liotta, T. Mchedlidze, G. Moroz, S. Whitesides, and S.xK. Wismath. Monotone simultaneous paths embeddings in R^{d}. Discrete Mathematics and Theoretical Computer Science, 20(1), January

2018. https://dmtcs.episciences.org/4177  .

[3]  L. Zeng, Shijie Xu, Y. Wang, K.xB. Kent,  , and C. Xu.

Toward cost-effective replica placements in cloud storage systems with QoS-awareness. Software: Practice and Experience, 47(6), 2017.

http://dx.doi.org/10.1002/spe.2441  .

[4] D. Bremner, T.xM. Chan, E.xD. Demaine, J. Erickson, F. Hurtado, J. Iacono, S. Langerman, M. Patrascu, and Perouz Taslakian. Necklaces, convolutions, and X + Y. Algorithmica, 69(2):294–314, December 2014. DOI:10.1007/s00453-012-9734-3.

http://arxiv.org/abs/1212.4771  .

[5] D. Bremner, M. Dutour, D.xV. Pasechnik, T. Rehn, and A. Schxrmann. Computing symmetry groups of polyhedra. LMS Journal of Computation and Mathematics, 17(1):565–581, 2014.

[6] D. Bremner, A. Deza, William Hua, and L. Schewe. More bounds on the diameters of convex polytopes. Optimization Methods and Software, 28(3):442–450, March 2013. DOI:10.1080/10556788.2012.668906B.

[7] Yoshitake Matsumoto, S. Moriyama, H. Imai, and D. Bremner. Matroid enumeration for incidence geometry. Discrete and Computational Geometry, 47(1):17-43, January 2012.

[8]  D. Bremner and L. Schewe.  Edge-graph diameter bounds

for convex polytopes with few facets. Experimental Mathematics, 20(3):229–237, October 2011. Funding: NSERC, (Hausdorff Institute),

AvH. http://arxiv.org/abs/0809.0915  .

[9]  D. Bremner, J. Bokowski, and G. Gevay.  Symmetric

matroid polytopes and their generation. European Journal of Combinatorics, 30:1758–1777, November 2009. Funding: NSERC.

http://www.cs.unb.ca/~bremner/research/papers/bbg-_smptg-_06.pdf  .

[10]  D.  Bremner,

M.xD. Sikirix, and A. Schxrmann. Polyhedral representation conversion up to symmetries. Polyhedral Computation, CRM Proceedings and Lecture Notes, 48:45–71, 2009. Funding: NSERC, (NB-Quebec).

http://www.arxiv.org/abs/math.MG/0702239  .

[11] D. Bremner, Dan Chen, J. Iacono, S. Langerman, and P. Morin. Output-sensitive algorithms for Tukey depth and related problems. Statistics and Computing, 18(3), September 2008. Funding: NSERC.

http://www.cs.unb.ca/~bremner/research/papers/bcilm-_osatd-_06.pdf  .

[12] O. Aichholzer, D. Bremner, E.xD. Demaine, F. Hurtado, E. Kranakis, H. Krasser, S. Ramaswami, S. Sethia, and J. Urritia. Games on triangulations. Theoretical Computer Science, 343(1-2):42–71, June 2005. Funding:

NSERC. http://dx.doi.org/10.1016/j.tcs.2005.05.007  .

[13] D. Bremner, E.xD. Demaine, J. Erickson, J. Iacono, S. Langerman, P. Morin, and G.xT. Toussaint. Output sensitive algorithms for computing nearest-neighbour decision boundaries. Discrete and Computational Geometry, 33(4):593–604, 2005. Funding: NSERC.

http://www.cs.unb.ca/~bremner/research/papers/bdeilmt-_osacn-_03b.ps.gz  .

[14]  D.  Bremner,  F.  Hurtado,  S.

Ramaswami, and V. Sacristxn. Small strictly convex quadrangulations of point sets. Algorithmica, 38(2):317–339, November 2003. Funding:

NSERC. http://www.arxiv.org/abs/cs.CG/0202011  .

[15] O. Aichholzer, D. Bremner, E.xD. Demaine, H. Meijer, V.

Sacristxn, and Michael Soss. Long proteins with unique optimal foldings in the H-P model. Computational Geometry: Theory and Applications, 25(1–2):139–159, May 2003. Funding: NSERC.

http://www.arXiv.org/abs/cs.CG/0201018  .

[16]  D. Bremner and T.xC. Shermer.  Point visibility graphs

and O-convex cover. International Journal of Computational Geometry and Applications, 10(1):55–71, February 2000.

http://www.cs.unb.ca/~bremner/research/papers/bs-_pvgoc-_00.ps.gz  .

[17] D. Bremner and V. Klee. Inner diagonals of convex polytopes. Journal of Combinatorial Theory A, 87(1):175–197, July 1999.

http://www.cs.unb.ca/~bremner/research/papers/bk-_idcp-_99.ps.gz  .

[18]  D.  Bremner.  Incremental  convex  hull

algorithms are not output sensitive. Discrete and

Computational  Geometry,  21(1):57–68,  January  1999.

http://www.cs.unb.ca/~bremner/research/papers/b-_ichan-_98.ps.gz  .

[19] M. dexBerg, P.xK. Bose, D. Bremner, S. Ramaswami, and G. Wilfong. Computing constrained minimum-width annuli of point sets. Computer-Aided Design, 30(4):267–275, April 1998.

http://dx.doi.org/10.1016/S0010-_4485(97)00073-_0  .

[20] P.xK. Bose, D. Bremner, and M. van Kreveld.

Determining  the  castability  of  simple  polyhedra.
Algorithmica,  19(1–2):84–113,  September  1997.

http://www.cs.unb.ca/~bremner/research/papers/bbk-_dcsp-_97.ps.gz  .

[21]  D. Avis, D. Bremner, and R. Seidel.  How good

are convex hull algorithms? Computational Geometry:

Theory  and  Applications,  7(5–6):265–301,  April  1997.

http://www.cs.unb.ca/~bremner/research/papers/abs-_hgach-_97.ps.gz  .

[22] D. Bremner. On the complexity of vertex and facet enumeration for convex polytopes. School of Computer Science, McGill University. 1997.

http://www.cs.unb.ca/~bremner/research/papers/phd  .

[23]  P.xK. Bose, D. Bremner, and G.xT. Toussaint.  All convex

polyhedra can be clamped with parallel jaw grippers. Computational Geometry: Theory and Applications, 6(5):291–302, September 1996.

http://www.cs.unb.ca/~bremner/research/papers/bbt-_acpcb-_96.ps.gz  .